# Is it possible to describe an electron by the evolution of a single point?

### v2

@@ 2.2 Dirac’s Hamiltonian
The last term is in fact ${\bi S}\cdot{\boldsymbol{\omega}}_{p}>0$, according to (9) and (13), because these two vectors are collinear but in opposite directions, as seen in figure 1. The spin $\bi S$} is orthogonal to the velocity $\bi u$}, and therefore to $\boldsymbol{\omega}_{u}$, and this term can be rewritten as ${\bi S}\cdot{\boldsymbol{\omega}}>0$. The energy contains two terms: the first term, proportional to the linear momentum, represents the translational energy, while the second, proportional to the spin and angular velocity and to the motion of the CC (zitterbewegung), and independent of the CM velocity, represents the rotational energy $H_{rot}$ which is positive definite, because the spin $\bi S$} has the opposite direction to the angular velocity $\boldsymbol{\omega}_{p}$, and never vanishes. If we introduce in (15) the expressions of $\bi p$} in (5) and $\bi S$} in (10),
@@ 2.3 Intrinsic properties of the Dirac particle
All physical observables $H$, $\bi p$}, the two spins $\bi S$}, ${\bi S}_{CM}$, the two components of the angular velocity $\boldsymbol{\omega}_{u}$ and $\boldsymbol{\omega}_{p}$ and the synchronous location of the CM $\bi q$} can be expressed in terms of the point $\bi r$} and their time derivatives. Also the electromagnetic properties magnetic moment $\boldsymbol{\mu}_{CM}$ and electric dipole moment ${\bi d}_{CM}$, if the particle has electric charge $e$.
In this analysis no hypothesis about the size or shape of the electron has been stated and no requirements concerning a possible charge or mass distribution. The physical properties of the electron are determined by the evolution of a single point, the center of charge. It is a system of only three degrees of freedom $\bi r$}, but they satisfy a system of ordinary differential equations of fourth order and not of second order, as it corresponds to the description of the spinless point electron. In three dimensional Euclidean space the most general continuous and differentiable curve satisfies necessarily a system of ordinary fourth order differential equations.
The historical suggestion in some physics books that the fundamental equations of physics are second order differential equations represents a simplifying restriction. This requires that a Lagrangian description of a moving point will be a function of the point and its velocity, Euler-Lagrange equations are of second order and thus the trajectories of the point have no torsion, which is equivalent to spin suppression. It represents the mechanical description of spinless particles but in the real world experimental physics has shown that there are no spinless elementary particles.
@@ 3 Natural units
The complete formalism can be done in terms of dimensionless variables once some basic properties are fixed. For the natural system of units we take as usual $\hbar=c=1$, the mass and charge of the electron $m=e=1$, the unit of length is $2R_{0}=\hbar/mc=1$, and the unit of time is $\tau_{0}=2R_{0}/c$, where $R_{0}$ is the separation between the CC and the CM for the center of mass observer. Then if we define new dimensionless position variables ${\widetilde{\bi r}}={\bi r}/2R_{0}$, ${\widetilde{\bi q}}={\bi q}/2R_{0}$, new dimensionless velocity variables ${\widetilde{\bi u}}={\bi u}/c$, ${\widetilde{\bi v}}={\bi v}/c$, and a new dimensionless time ${\widetilde{t}}={t}/\tau_{0}$, the differential equations are written in terms of dimensionless variables, and once the tildes are removed, they become:
@@
Therefore this system (19) and (20) of ordinary non-linear differential equations, written in natural units, depend on two physical dimensionless coefficients, $K_{E}E$ and $K_{B}B$, where $E$ and $B$ are, the intensity of the electric and magnetic field, respectively, in the I.S. of units, and all remaining variables are dimensionless. The stability of this non-linear system of ordinary differential equations depends on the values of these two coefficients and therefore sets a limit on the allowed values of the external fields.
The relationship between the International System of units and the Natural System of units has the following conversion factors:
@@ 5 Mass and Spin invariance: Radiation reaction
By energy conservation if this difference of energy ${\bi F}\cdot(d{\bi r}-d{\bi q})$, is positive the electron increases its mechanical energy, and the remaining part of this energy that has not been used in this acceleration process, is transformed into energy of the field. If this difference is negative, the electron loses mechanical energy and the difference is returned and transformed into energy of the field. It can be interpreted as if the particle radiates electromagnetic energy, because the total electromagnetic energy of the field has changed. If these two works are the same there is no radiation because all electromagnetic energy of this work is transformed into mechanical energy of the particle.
@@
The variation of the linear momentum contains a breaking term, opposite to the CM velocity $\bi v$}, that depends on the factor $\gamma(v)^{2}$ and the difference of the work per unit time of the external Lorentz force along the CC and the CM trajectories. This term can be interpreted as the reaction force produced by the interchange of photons between the particle and the field. In a continuous interaction, the term $\gamma(v)^{2}\left[{\bi F}\cdot({\bi u}-{\bi v})\right]d{\bi q}$, will be the linear momentum of the continuously radiated electromagnetic field in a time $dt$.
@@
A single electron in a finite time $T$ radiates finite electromagnetic energy. This corresponds to the emission of a finite number of photons, let us say $N$. This means that when a single electron is accelerated it is not emitting continuous radiation but rather it emits on average a single photon after a discrete time $\Delta t=T/N$. Radiation is not a continuous process, so that we need to wait to a discrete time $\Delta t$ such that the quantum of energy corresponds to the transfer of angular momentum $\hbar$. The determination of this discrete time in different examples is left to a future research.
@@ 6 Conclusions
Radiation is not a continuous process. Nevertheless the classical description of the electron and Maxwell’s equations are based on continuous and differentiable variables and fields. If according to Planck’s hypothesis radiation is the emission of photons, these are not emitted continuously, and a discrete time must exist between two consecutive emission of photons by any single electron, until the continuously radiated electromagnetic angular momentum amounts the quantum $\hbar$.
The complete set of dynamical equations can be expressed in terms of dimensionless variables or natural units, where the only physical parameters that characterize the system of equations are the intensity of the electric and magnetic fields in natural units, so that the stability of these equations will be restricted by these values. The allowed electric and magnetic fields are not unbounded. The existence of this discrete time $\Delta t$ between the emission of consecutive photons implies that the allowed values of the electric and magnetic field are bounded above if the dynamical equations have to remain stable. For greater values of the fields the constraint that the CM velocity $v<1$, does not hold.

### v3

@@ 2.2 Dirac’s Hamiltonian
The last term is always positive definite ${\bi S}\cdot{\boldsymbol{\omega}}_{p}>0$, according to (9) and (13), because these two vectors are collinear but in opposite directions, as seen in figure 1. The spin $\bi S$} is orthogonal to the velocity $\bi u$}, and therefore to $\boldsymbol{\omega}_{u}$, and this term can be rewritten as ${\bi S}\cdot{\boldsymbol{\omega}}>0$. The energy contains two terms: the first term, proportional to the linear momentum, represents the translational energy, while the second, proportional to the spin and angular velocity and to the motion of the CC (zitterbewegung), and independent of the CM velocity, represents the rotational energy $H_{rot}$ which is positive definite, because the spin $\bi S$} has the opposite direction to the angular velocity $\boldsymbol{\omega}_{p}$, and never vanishes. If we introduce in (15) the expressions of $\bi p$} in (5) and $\bi S$} in (10),
@@ 2.3 Intrinsic properties of the Dirac particle
All physical observables $H$, $\bi p$}, the two spins $\bi S$}, ${\bi S}_{CM}$, the two components of the angular velocity $\boldsymbol{\omega}_{u}$ and $\boldsymbol{\omega}_{p}$ and the synchronous location of the CM $\bi q$} can be expressed in terms of the point $\bi r$} and their time derivatives, simultaneously. Also, the electromagnetic properties magnetic moment $\boldsymbol{\mu}_{CM}$ and electric dipole moment ${\bi d}_{CM}$, if the particle has electric charge $e$.
In this analysis no hypothesis about the size or shape of the electron has been stated and no requirements concerning a possible charge or mass distribution. The physical properties of the electron are determined by the evolution of a single point, the center of charge. It is a system of only three degrees of freedom $\bi r$}, but they satisfy a system of ordinary differential equations of fourth order and not of second order, as it corresponds to the description of the spinless point electron. In three-dimensional Euclidean space the most general continuous and differentiable curve satisfies necessarily a system of ordinary fourth order differential equations.
The historical suggestion in some physics books that the fundamental equations of physics are second order differential equations represents a simplifying restriction. This requires that a Lagrangian description of a moving point will be a function of the point and its velocity, Euler-Lagrange equations are of second order and thus the trajectories of the point have no torsion, which is equivalent to spin suppression. It represents the mechanical description of spinless particles but in the real-world experimental physics has shown that there are no spinless elementary particles.
@@ 3 Natural units
The complete formalism can be done in terms of dimensionless variables once some basic properties are fixed. For the natural system of units we take as usual $\hbar=c=1$, the mass and charge of the electron $m=e=1$, the unit of length is $2R_{0}=\hbar/mc=1$, and the unit of time is $\tau_{0}=2R_{0}/c=1$, where $R_{0}$ is the separation between the CC and the CM for the center of mass observer. Then if we define new dimensionless position variables ${\widetilde{\bi r}}={\bi r}/2R_{0}$, ${\widetilde{\bi q}}={\bi q}/2R_{0}$, new dimensionless velocity variables ${\widetilde{\bi u}}={\bi u}/c$, ${\widetilde{\bi v}}={\bi v}/c$, and a new dimensionless time ${\widetilde{t}}={t}/\tau_{0}$, the differential equations are written in terms of dimensionless variables, and once the tildes are removed, they become:
@@
Therefore, this system (19) and (20) of ordinary non-linear differential equations, written in natural units, depend on two physical dimensionless coefficients, $K_{E}E$ and $K_{B}B$, where $E$ and $B$ are, the intensity of the electric and magnetic field, respectively, in the I.S. of units, and all remaining variables are dimensionless. The stability of this non-linear system of ordinary differential equations depends on the values of these two coefficients and therefore sets a limit on the allowed values of the external fields.
[…]
@@ 5 Mass and Spin invariance: Radiation reaction
By energy conservation if this difference of energy ${\bi F}\cdot(d{\bi r}-d{\bi q})$, is positive the electron increases its mechanical energy, and the remaining part of this energy that has not been used in this acceleration process, must remain as energy of the field. If this difference is negative, the electron loses mechanical energy and the difference is returned and transformed into energy of the field. It can be interpreted as if the particle radiates electromagnetic energy, because the total electromagnetic energy of the field has changed. If these two works are the same there is no radiation because all electromagnetic energy of this work is transformed into mechanical energy of the particle.
@@
The variation of the linear momentum contains a breaking term, opposite to the CM velocity $\bi v$}, that depends on the factor $\gamma(v)^{2}$ and the difference of the work per unit time of the external Lorentz force along the CC and the CM trajectories. This term can be interpreted as the reaction force produced by the interchange of photons between the particle and the field. In a continuous interaction, the term $\gamma(v)^{2}\left[{\bi F}\cdot({\bi u}-{\bi v})\right]d{\bi q}$, will be the linear momentum of the continuously radiated electromagnetic field in a time $dt$, that is substracted from the impulse ${\bi F}dt$ produced by the external force.
@@
A single electron in a finite time $T$ radiates finite electromagnetic energy. This corresponds to the emission of a finite number of photons, let us say $N$. This means that when a single electron is accelerated it is not emitting continuous radiation but rather it emits on average a single photon after a discrete time $\Delta t=T/N$. Radiation is not a continuous process, so that we need to wait to a discrete time $\Delta t$ such that the quantum of energy corresponds to the transfer of angular momentum $\hbar$. The determination of this discrete time between two consecutive emissions of photons by a single electron in different examples, is left to future research.
6 Fine structure constant
In (2007) we have found a Poincaré invariant interaction Lagrangian between two Dirac particles [12]. If $x_{a}^{\mu}\equiv(ct_{a},{\bi r}_{a})$, $a=1,2$ are the 2 four-vectors of both particles and $\dot{x}_{a}^{\mu}\equiv(c\dot{t}_{a},\dot{\bi r}_{a})$, $a=1,2$, are the $\tau$-derivatives with respect to some invariant evolution parameter $\tau$, they transform as $x_{a}^{{}^{\prime}\mu}=\Lambda^{\mu}_{\nu}x_{a}^{\nu}+a^{\mu}$, and $\dot{x}_{a}^{{}^{\prime}\mu}=\Lambda^{\mu}_{\nu}\dot{x}_{a}^{\nu}$, where $\Lambda^{\mu}_{\nu}$ is a general Lorentz transformation and $a^{\mu}$ a space-time translation. The following Poincaré invariant interaction Lagrangian
$\widetilde{L}_{I}=g\sqrt{\frac{\dot{x}_{1}^{\mu}\dot{x}_{2\mu}}{(x_{1}-x_{2})^{\mu}(x_{2}-x_{1})_{\mu}}}=g\sqrt{\frac{c^{2}\dot{t}_{1}\dot{t}_{2}-\dot{\bi r}_{1}\cdot\dot{\bi r}_{2}}{({\bi r}_{1}-{\bi r}_{2})^{2}-c^{2}(t_{1}-t_{2})^{2}}},$
depends on the coupling constant $g$ with dimensions of action if the evolution parameter $\tau$ is taken dimensionless. This interaction Lagrangian is also space-time scale invariant under a scale transformation that preserves the speed of light and also symmetric in the interchange $1\leftrightarrow 2$. When selecting some inertial observer, the evolution parameter is taken the time coordinate $\tau=t$ in this frame, and the evolution description is a synchronous evolution, and thus $t_{1}=t_{2}=t$, $\dot{t}_{a}=1$, and the synchronous Lagrangian in this frame becomes
$L_{I}=g\sqrt{\frac{c^{2}-{\bi u}_{1}\cdot{\bi u}_{2}}{({\bi r}_{1}-{\bi r}_{2})^{2}}}.$
This Lagrangian, without tilde in the time evolution description, has now dimensions of energy, is translation and rotation invariant but not invariant under pure Lorentz transformations, because the two particles are considered with simultaneity in this frame, and simultaneous events in one frame are not simultaneous in another moving frame.
With that Lagrangian we have analyzed numerically [20] different electron-electron interactions. If we introduce the extra assumption that when the spin vanishes the CC and the CM become the same point, ${\bi u}_{1}={\bi u}_{2}=0$, the above interaction Lagrangian between the two Dirac particles becomes the instantaneous Coulomb interaction Lagrangian between two spinless point particles at rest, the coupling constant between the two Dirac particles is just $g=e^{2}/4\pi\epsilon_{0}c$. This Lagrangian has dimensions of energy. The transformation coefficient from the International System of units to the Natural System for energy is the factor $1/mc^{2}$. The above interaction Lagrangian when written in natural units, and including the energy conversion factor, becomes:
$L_{I}=\frac{1}{mc^{2}}\frac{e^{2}}{4\pi\epsilon_{0}c}\frac{c}{2R_{0}}\sqrt{\frac{1-{\bi u}_{1}\cdot{\bi u}_{2}}{({\bi r}_{1}-{\bi r}_{2})^{2}}}=\alpha\sqrt{\frac{1-{\bi u}_{1}\cdot{\bi u}_{2}}{({\bi r}_{1}-{\bi r}_{2})^{2}}},\quad\alpha=\frac{e^{2}}{4\pi\epsilon_{0}\hbar c},$
$\alpha\approx 1/137$, with all variables in natural units and $u_{1}=u_{2}=1$.
It is only through this non-relativistic limit of a Poincaré-invariant interaction Lagrangian and that the basic instantaneous interaction between point particles is given by Coulomb’s law, that we have arrived at the conclusion that the only dimensionless parameter characterizing the interaction between two Dirac particles is the fine-structure constant. The existence of this dimensionless constant implies that the fundamental constants $\hbar$, $e$, $c$ and the permittivity of vacuum $\epsilon_{0}$, are not independent physical parameters.
7 Gravitational interaction
The above non-relativistic limit of a Poincaré invariant Lagrangian suggests to consider whether we can arrive similarly to the non-relativistic Newtonian gravitational potential between two Dirac particles, as a consequence of their gravitational interaction.
The theory of gravitation is today Einstein’s General Theory of Relativity. In the spirit of unification of all interactions, is it possible to find a Lagrangian description of the gravitational interaction between two spinning Dirac particles? In this case the coupling constant of the gravitational interaction of two Dirac particles is known as the gravitational coupling constant, and we get:
$\alpha_{G}=\frac{Gm^{2}}{\hbar c}=1.75175\cdot 10^{-45},$
where $G$ is the gravitational constant and $m$ is the mass of the electron.
But this analysis of the gravitational interaction of two electrons has led us to the conclusion that today’s gravitational theory is a very restrictive theory of gravitation.
Einstein’s theory was formulated before the discovery in the formalism of the calculus of variations that the manifold $X$ of the boundary variables manifold of the Lagrangian description of any arbitrary mechanical system is always a metric space. It is this manifold that in this approach supplies the possible degrees of freedom and essential variables of an elementary particle.
Euler-Lagrange equations are equivalent to the geodesic equations on that manifold between the initial point $x_{1}\in X$ to the final point $x_{2}\in X$, such that the distance along the path is a minimum. This discovery is the result of the work of Finsler (1918) [13], followed later by the paper by Cartan (1933) who called Finsler spaces to these geometric structures [14]. More detailed books are those of Rund (1959) [15] and Asanov (1985) [16].
The metric structure of the boundary variables manifold is characterized by a symmetric tensor $g_{ij}(x,\dot{x})$ which is a function of the point $x\in X$ but also of the derivative $\dot{x}$ with respect to some arbitrary and invariant evolution parameter $\tau$. The reason is that the Lagrangian, if the evolution is analyzed in terms of some arbitrary and invariant evolution parameter $\tau$, is a function $\widetilde{L}(x,\dot{x})$ of the boundary variables and their first order derivative with respect to $\tau$. But it is a homogeneous function of first degree in terms of the derivatives $\dot{x}_{i}$ of all the boundary variables.
The action integral is usually stated in terms of a time integral. But time is a relative observable and therefore the formalism expressed in terms of some particular time is not independent of the inertial observers. If we make the action integral in terms of some arbitrary invariant evolution parameter $\tau$, the same for all inertial observers, can be rewritten as:
$\int_{t_{1}}^{t_{2}}L\,dt=\int_{\tau_{1}}^{\tau_{2}}L\,\dot{t}(\tau)d\tau=\int_{\tau_{1}}^{\tau_{2}}\widetilde{L}\,d\tau,\quad\widetilde{L}=L\dot{t}(\tau),$
and $\widetilde{L}$, is a homogeneous function of first degree in the $\tau$-derivatives of the boundary variables [6]. If the evolution parameter $\tau$ is dimensionless, this $\widetilde{L}$ has dimensions of action.
This homogeneity allows us to write the Lagrangian of any mechanical system with $n$ boundary variables $x_{i}$, $i=1,\ldots,n$, as a sum of as many terms as boundary variables in the form
$\widetilde{L}=\sum_{i=1}^{i=n}\frac{\partial\widetilde{L}}{\partial\dot{x}_{i}}\dot{x}_{i}=\sum_{i}F_{i}(x,\dot{x})\dot{x}_{i},$
where the $F_{i}$ are homogeneous functions of zero degree of the $\dot{x}_{i}$, and thus depend on the time derivatives of the independent degrees of freedom.
The Lagrangian squared $\widetilde{L}^{2}$ is a homogeneous function of second degree of all derivatives $\dot{x}_{i}$. This allows us to write it as
$\widetilde{L}^{2}=\sum_{ij}g_{ij}(x,\dot{x})\dot{x}_{i}\dot{x_{j}},\quad g_{ij}(x,\dot{x})=\frac{1}{2}\frac{\partial^{2}\widetilde{L}^{2}}{\partial\dot{x}_{i}\partial\dot{x}_{j}}=g_{ji},$
and the partial derivatives of $\widetilde{L}^{2}$ define the components of the symmetric metric tensor. The action integral can be rewritten as the integral of a positive function
$\int_{\tau_{1}}^{\tau_{2}}\widetilde{L}(x,\dot{x})d\tau=\int_{\tau_{1}}^{\tau_{2}}\sqrt{\widetilde{L}^{2}(x,\dot{x})}d\tau=\int_{\tau_{1}}^{\tau_{2}}\sqrt{g_{ij}(x,\dot{x})\dot{x}^{i}\dot{x}^{j}}d\tau=$
$\int_{x_{1}}^{x_{2}}\sqrt{g_{ij}(x,\dot{x})d{x}^{i}d{x}^{j}}=\int_{x_{1}}^{x_{2}}ds,$
and the action integral is the positive geodesic distance between the extremal points.
As an example, the free relativistic point particle Lagrangian
$L_{0}(t,{\bi r},{\bi u})=-mc\sqrt{c^{2}-{\bi u}^{2}},\quad\widetilde{L}_{0}(t,{\bi r},\dot{t},\dot{\bi r})=-mc\sqrt{c^{2}\dot{t}^{2}-\dot{\bi r}^{2}},$
$\widetilde{L}_{0}$ is a squared root of a second-degree polynomial in terms of all the $\dot{x}^{\mu}$ and therefore a homogeneous function of first degree in these derivatives. In natural units
$\frac{1}{2}\frac{\partial^{2}\widetilde{L}_{0}^{2}}{\partial\dot{x}^{\mu}\partial\dot{x}^{\nu}}=\eta_{\mu\nu},\quad\eta_{\mu\nu}={\rm diag}(1,-1,-1,-1),$
where $\eta_{\mu\nu}$ is Minkowski’s metric tensor. The metric of the boundary variables manifold of the free relativistic point particle is Minkowski’s metric. It is a constant Finsler metric, independent of the point and of the velocity. If this particle interacts and $\widetilde{L}_{I}$ is the interaction Lagrangian, the new metric under interaction of the same manifold is obtained by derivation of the function $(\widetilde{L}_{0}+\widetilde{L}_{I})^{2}$. But this new metric depends of the point $x^{\mu}$ and also on the $\dot{x}^{\mu}$ and it is not a Riemannian metric. Any interaction modifies the Finsler metric of the point particle.
General relativity postulates that gravity modifies the metric of the space-time or, concerning the test particle, modifies the metric of its boundary variables manifold. That postulate is unnecessary because all interactions modify the metric. The statement that the modified metric is a new tensor $g_{\mu\nu}(x)$, is a very strong restriction because the modification of the metric should include dependence also on the derivatives $g_{\mu\nu}(x,\dot{x})$. The assumption that the modified metric tensor is a pseudo-Riemannian metric is a very strong constraint because the variational formalism implies that the new modified metric is a Finsler metric. If the velocity is negligible with respect to the speed of light, implies that in the low velocity limit
$\lim_{v/c\to 0}g_{\mu\nu}(x,\dot{x})=g_{\mu\nu}(x),$
produces a Riemannian metric.
We have analyzed this constraint [17] and shown that General Relativity is a low velocity limit of a more general metric description of space-time. Application of General Relativity to the solar system produces reasonable results because the Sun and planets move with a small velocity compared with $c$. High velocity gravitational phenomena like very high red-shifts or gravitational collapse, which could be dependent of the velocity are left outside this possible description because Riemannian metrics are velocity independent. Other conclusions of this very restrictive theory of gravitation have to be taken with caution, including the singularities of the Schwarzschild metric.
The general solution of Einstein’s equations can never produce a Finsler metric, because the formalism is organized to obtain as solutions only Riemannian metrics. The fundamentals of General Relativity have to be revisited to allow Finsler metric structures as solutions of the general gravitational equations.
@@ 8 Conclusions
Radiation is not a continuous process. Nevertheless, the classical description of the electron and Maxwell’s equations are based on continuous and differentiable variables and fields. If according to Planck’s hypothesis radiation is the emission of photons, these are not emitted continuously, and a discrete time must exist between two consecutive emissions of photons by any single electron, until the continuously radiated electromagnetic angular momentum amounts the quantum $\hbar$.
The complete set of dynamical equations can be expressed in terms of dimensionless variables or natural units. In the electromagnetic interaction the only physical parameters that characterize the system of equations are the intensity of the electric and magnetic fields in natural units, so that the stability of these equations will be restricted by these values. The allowed electric and magnetic fields are not unbounded. The existence of this discrete time $\Delta t$ between the emission of consecutive photons implies that the allowed values of the external electric and magnetic fields are bounded above if the dynamical equations have to remain stable. For greater values of the fields the constraint that the CM velocity $v<1$, does not hold.
The fine structure constant is the only dimensionless parameter that characterizes a Poincaré invariant interaction between two Dirac particles, if the non-relativistic limit of this interaction is the Coulomb interaction.
The possible gravitational interaction between two electrons has led us to the conclusion that the foundations of General Relativity have to be revised in order to obtain Finsler metrics rather than Riemannian metrics for the description of gravitational phenomena.